# CollabStory: Multi-LLM Collaborative Story Generation and Authorship Analysis

### v1

@@ 4.2 Task 2: How many authors have written a story?
The second task is to predict the number of authors involved in generating a story. For the CollabStory dataset, this means that class labels $\in[1,5]$. From Table 7 we see that the task of predicting exactly how many authors have co-written a story is easiest for $N=1$ in conjunction with findings from Task 1 that showed that multi-authored text can be more easily distinguished from single-authored text. Thus, here too it seems to be easiest to separate the single-authored texts from $N\geq 2$. However, for multi-authored stories, only BERT and RoBERTa perform better than other baselines (>0.72 F1), especially for $N\in[4,5]$. Overall, the performance across this task is low.
Task 2: Predict Number of Authors (N)
Method N=1 N=2 N=3 N=4 N=5
MNB 0.72 0.50 0.49 0.48 0.51
SVM 0.68 0.43 0.52 0.40 0.54
BERT 0.79 0.70 0.64 0.75 0.81
ALBERT 0.70 0.57 0.55 0.53 0.65
RoBERTa 0.76 0.68 0.63 0.72 0.74
AVG 0.73 0.58 0.57 0.58 0.65
Table 7: All scores are F1-scores. AVG denotes average F1-score for each $N$. This task is particularly challenging with only the single-authored stories ($N=1$) being correctly classified. For all multi-authored texts, BERT and RoBERTa perform better than others. Best performing method is in bold and second highest underlined for each N.
@@ 4.3 Task 3: Authorship Verification
This is a pair-wise sentence classification task where the goal is to predict if two adjacent sentences are written by the same author or not. For this task, we used all the sentences at LLM-LLM boundaries, that is the last sentence of part $i$ and the first sentence of part $i+1$. The negative class data samples were sampled as random pairs of consecutive sentences within each story part.
Task 3: Authorship Verification
Method N=2 N=3 N=4 N=5
MNB 0.65 0.64 0.60 0.63
SVM 0.63 0.62 0.62 0.63
BERT 0.73 0.73 0.73 0.71
ALBERT 0.92 0.89 0.89 0.89
RoBERTa 0.91 0.91 0.90 0.89
AVG 0.76 0.75 0.74 0.75
Table 8: All scores are F1-scores and AVG denotes average F-1 scores for each $N$ for the task of detecting authorship boundaries between sentence pairs. We see that generally, performance is slightly higher for $N=2$ across all classifiers. Best performing method is in bold and second highest underlined for each N.
From Table 8, we see that transformers-based fine-tuned methods perform well at this task. We also note that detecting sentence authorship boundaries seems to be slightly easier for the 2-author case than for $N\geq 3$.
@@ 4.4 Task 4: Authorship Attribution
Authorship Attribution involves predicting exactly who the author of a text article is. In the case of multi-LLM text, we design this task such that each data sample is homogeneous or each part is written by a single author and the classifier’s task is to identify its author.
@@
From Table 9, we see that most of the authors seem hard to identify irrespective of the value of $N$, except for Gemma. We were expecting attribution to be easier the fewer the number of authors since the length of the parts contributed by each author would be longer. But there does not seem to be any such correlation in our dataset i.e. length contributed by each author does not correlate with their detection. To further explore why Gemma was easily identifiable, we manually inspected a random sample of 100 articles for discernible features or peculiarities in story parts written by Gemma v/s all other authors. We provide examples of story parts generated by Gemma and other LLMs in the Appendix (Table 11). We found that the text generated by Gemma is not noticeably different from that of other authors. We leave a deeper analysis of potential factors, such as training data-induced biases, story part, or author order for future work.
5 Discussion
Recent developments have significantly advanced LLM-assisted writing, sparking widespread discussions about the nature of authorship. Beyond using LLMs for paraphrasing, editing, and enhancing text, there exists an extreme scenario where text is generated entirely by multiple LLMs. Our work addresses this extreme case, raising several nuanced authorship concerns: Who should be considered the true creative source in such a situation? Should all LLMs involved be credited? Or should the human developers designing the prompts be acknowledged as the primary authors? Moreover, should the LLM that contributed the most—whether in terms of word count, narrative depth, or plot twists—be granted greater ownership? There have been recent works addressing this question of ownership. For example, Joshi and Vogel (2024) and Dhillon et al. (2024) found that humans felt a sense of higher ownership when they wrote longer prompts and when the AI assistant-generated text length was shorter, respectively. The question of authorship when text is repeatedly paraphrased using LLMs has also been deeply considered by Tripto et al. (2023), who find that the notion of authorship is task-dependent and cannot be generalized.
These questions have profound implications for various stakeholders in the burgeoning socio-technical system of generative AI. Our research introduces authorship-related tasks using CollabStory, which can help address these concerns by accurately discerning the usage of multiple LLMs in texts. Our extension of PAN-inspired authorship tasks is closely linked to real-world implications, as follows:
@@ 5 Discussion
6 Conclusion
In this work, we present CollabStory, the first exclusively LLM-LLM or machine-machine collaborative story dataset, and demonstrate the tasks it enables. We elaborate on why multi-LLM authorship tasks are crucial to study by discussing their real-world implications (summarized in Table 10). We will soon need “Catch As Catch Can” methods to not only find all points where authorship changes within an article (style change detection) but also simultaneously attribute each independent segment to the specific LLM author (attribution). As more and more LLMs are becoming easier to access, malicious actors could combine texts from different LLMs to evade automated and in-built misinformation flaggers, or students might circumvent credibility checks by having different LLMs write different sections of an academic article. Thus, CollabStory was developed as a resource with long-form stories written by multiple LLMs to support the development and expansion of tasks and methods that can help address incoming challenges brought by LLM-LLM interactions.
Limitations
Ethics Statement
Appendix A Appendix
A.1 Examples of story parts written by Gemma followed by other authors
We inspected 100 randomly sampled stories and inspected the parts written by Gemma in search of visibly discernible features or peculiarities that might have explained the ease of its detection as compared to other LLMs. To the naked eye, this text seemed to be similar to all the other parts of the story since we subjected all the story parts to the same filtration process. We suspect that other factors might be in play here such as the author order or tone differences, and leave this exploration to future attempts (Table 11 with examples of story samples in next page).
A.2 LLM Prompting
A.3 Dataset Cleaning
Gemma Not Gemma
The chase has begun, and adrenaline floods his veins. The blinding headlights pierce through the night, illuminating the frantic dance of tires on asphalt. The adrenaline-fueled chase weaves through the labyrinth of city streets, sirens screaming like tormented spirits. The driver, hands white on the wheel, desperately maneuvers, adrenaline masking the fear gnawing at his conscience. Every corner, every alleyway, feels like a potential trap, each sudden movement a potential assault. The police officer, unwavering in his pursuit, his seasoned instincts guiding him through the chaos. He maneuvers his cruiser with precision, anticipating the driver’s next move. His heart pounds in his chest, adrenaline battling with the growing sense of urgency. He knows the city like the back of his hand, weaving through traffic, his eyes glued to the fleeing vehicle. As the chase intensifies, the driver’s desperation escalates. He risks collisions, weaving through pedestrians, pushing the limits of his vehicle’s performance. The officer remains calm, focused, his training taking over. He knows he must bring this chase to an end, but he also understands the risks involved. Swallowing his fear, he decides to take the next exit and lead his pursuers through the winding, narrow streets of the old town. The streets are alive with the sound of market vendors calling out their wares, children laughing, and the faint sounds of a distant band playing music. The narrow, cobblestone alleys are lined with ancient buildings, each one more beautiful than the last. He knows that this labyrinth of streets will give him an advantage, but it also increases the risk of colliding with a pedestrian or another car. Despite the danger, he continues to navigate the maze of alleys, turning sharply at each intersection. He can hear the revving of engines and the screeching of tires behind him, but he refuses to look back. He knows that every twist and turn could be his last, but he is determined to outsmart his pursuers and escape. As he rounds the final corner, he sees the safety of the forest beyond the old town walls. He grits his teeth, puts the pedal to the metal, and bursts through the gates, leaving the chaos of the chase behind him.
Emily stood frozen, her heart pounding in her chest. A sense of awe washed over her as she approached the base of the tree and noticed a faint glow emanating from beneath the ground. Curiosity overwhelmed her, and she cautiously lowered a hand into the soft earth. As her fingers brushed against something cold and smooth, a surge of energy flowed through her body. A deep voice rumbled in her mind, "You have awakened. The spirit of the wolf within you has been awakened." The earth trembled beneath her feet, and a swirling mist began to envelop Emily. She felt her bones shifting, her muscles growing taut. Her reflection in the nearby pond transformed from a girl to a magnificent silver wolf. She barked, a primal call echoing through the woods. The glow from beneath the oak tree intensified, and a majestic wolf spirit emerged, its eyes glowing with wisdom and power. The spirit guided Emily through the forest, teaching her the secrets of her newfound lineage. She learned of the ancient bond between wolves and humans, the importance of loyalty and bravery. As she embraced her new identity, she found herself a part of an intricate web of relationships. Her new pack, made up of both humans and wolves, accepted her with open arms. Together, they faced challenges, both natural and supernatural. Emily’s courage and quick thinking saved the pack on more than one occasion. Her transformation brought her closer to her true self, and she no longer felt lost or alone. The bonds of loyalty and friendship grew stronger as they worked together to protect their territory. Emily felt a deep sense of belonging, and she knew that she had found her place in the world. The woods, once a place of solitude, had become her home, a place of love and acceptance. Emily’s life as a wolf shifter was not without its difficulties, but she faced each challenge with the same bravery and determination that had always defined her. She had found her true identity, and nothing would ever shake her sense of belonging or her commitment to her pack. As the sun set over the glade, Emily looked up at the stars, her heart filled with gratitude and love.
In the unspoken words hanging between us, Sarah and I understand that the journey to healing will be arduous. There will be moments of doubt, days when the weight feels unbearable, when the waves of sadness threaten to pull us under. But we are not in this together. We are stronger than our individual struggles. Together, we will learn to navigate the storms, to embrace the sunshine, to rediscover who we are beyond the baggage we once carried. Sarah’s smile brightens my day, and I know that this new chapter in our lives is going to be filled with adventure, growth, and most importantly, love. We’ll stumble, we’ll trip, but we’ll always have each other to pick us back up and move forward, stronger than before. I take Sarah’s hand and squeeze it tightly.
Her words seemed to hang in the air like acrid smoke, but I knew better. The voice, the chilling certainty in its delivery, had belonged to something far more sinister. I glanced at the others, their faces pale, their eyes wide with disbelief. "There’s no way this is real," one whispered, her voice trembling. But the undeniable presence of death in our living room told us otherwise. The figure shifted, its skeletal hands brushing the floor with an unholy whisper. "Your time has come," it hissed, its voice like nails on a chalkboard. Panic surged through me, my limbs frozen in terror. The others stumbled around, scrambling for escape, but their movements were clumsy, their screams muted by fear. In that moment, I realized that the only way out was through the front door, directly into the path of the creature that had come for me. With a surge of adrenaline, I broke free from the group, sprinting towards the exit. As I reached the door, it swung open to reveal Tia, the figure from my nightmares, standing there. But instead of fear, I felt relief. She had a warm smile on her face, and her eyes were full of compassion. "You have been chosen to bear the mark of the guardian," she explained, holding out the hourglass. "I have protected this town for generations, and now it’s your turn to carry on the tradition." I hesitated for a moment, but then I accepted the hourglass with a sense of pride and purpose. Together, Tia and I faced the darkness that lurked outside, ready to protect our town and its people from the evil that sought to harm them. From that day on, I became the guardian of my town, using the powers granted to me by Tia to keep the peace and protect those around me. The people of the town were grateful, and I knew that I had found a new purpose in life.
Table 11: As can be seen from a few examples, the text written by Gemma is not necessarily different from that of other authors (in the second column). Thus, from our manual inspection, we were unable to attribute any noticeable features to the higher authorship attribution performance for Gemma.
Feature Author K=1 K=2 K=3 K=4 K=5
Llama 0.60 ± 0.05 0.61 ± 0.05 0.59 ± 0.05 0.58 ± 0.05 0.58 ± 0.06
mistral 0.64 ± 0.04 0.62 ± 0.04 0.61 ± 0.05 0.61 ± 0.04 0.61 ± 0.05
Olmo 0.62 ± 0.05 0.57 ± 0.06 0.58 ± 0.06 0.58 ± 0.06 0.57 ± 0.07
Orca 0.62 ± 0.05 0.62 ± 0.04 0.61 ± 0.04 0.61 ± 0.05 0.60 ± 0.05
Llama 83.11 ± 8.35 82.61 ± 8.50 82.75 ± 9.00 80.88 ± 10.10 80.13 ± 9.19
Mistral 81.04 ± 8.58 83.99 ± 8.40 81.59 ± 10.07 82.02 ± 9.09 79.91 ± 9.44
Olmo 80.78 ± 9.01 83.31 ± 9.87 81.41 ± 9.81 80.55 ± 10.73 80.45 ± 10.20
Orca 83.08 ± 8.54 82.51 ± 8.45 80.86 ± 9.65 79.95 ± 9.97 79.51 ± 9.80
Coherence Gemma 0.49 ± 0.07 0.47 ± 0.08 0.47 ± 0.08 0.47 ± 0.08 0.48 ± 0.08
Llama 0.44 ± 0.08 0.47 ± 0.08 0.47 ± 0.08 0.47 ± 0.08 0.49 ± 0.08
mistral 0.46 ± 0.07 0.44 ± 0.07 0.45 ± 0.07 0.45 ± 0.07 0.46 ± 0.07
Olmo 0.47 ± 0.08 0.43 ± 0.07 0.45 ± 0.08 0.45 ± 0.08 0.45 ± 0.08
Orca 0.44 ± 0.07 0.45 ± 0.08 0.45 ± 0.08 0.46 ± 0.08 0.47 ± 0.07

### v2

@@ 5.2 Task 2: How many authors have written a story?
The second task is to predict the number of authors involved in generating a story. For the CollabStory dataset, this means that class labels $\in[1,5]$. From Figure 3(b) we see that the task of predicting exactly how many authors have co-written a story is easiest for $N=1$ in conjunction with findings from Task 1 that showed that multi-authored text can be more easily distinguished from single-authored text. Thus, here too it seems to be easiest to separate the single-authored texts from $N\geq 2$. However, for multi-authored stories, only BERT and RoBERTa perform better than other baselines (>0.72 F1), especially for $N\in[4,5]$. Overall, the performance across this task is low.
@@ 5.3 Task 3: Authorship Verification
This is a pair-wise sentence classification task where the goal is to predict if two consecutive sentences are written by the same author or not. For this task, we used all the sentences at LLM-LLM boundaries, that is the last sentence of part $i$ and the first sentence of part $i+1$. The negative class data samples were sampled as random pairs of consecutive sentences within each story part. From Figure 3(c), we see that transformers-based fine-tuned methods perform well at this task (> 0.8 F1).
@@ 5.4 Task 4: Authorship Attribution
Authorship Attribution involves predicting exactly who the author of a text article is. In the case of multi-LLM text, we design this task such that each data sample is homogeneous or each part is written by a single author and the classifier’s task is to identify its author. From Figure 3(d), we see that irrespective of the value of $N$, attribution performance is low (between 0.6 to 0.75 F1). We were expecting attribution to be easier the fewer the number of authors in the article since the length of the parts contributed by each author would be longer. But there does not seem to be any such correlation in our dataset i.e. length contributed by each author does not correlate with their detection. We provide more detailed results for each LLM or class-wise F1 scores in the Appendix (Table 8).
Thus, from the perspective of authorship tasks, for the multi-LLM scenario, Task 1 and Task 3 have higher performance (>=0.8 F1) and it is relatively easier to predict if a text is single or multi-authored and if two consecutive texts are written by the same author. , Task 2 and Task 4 are challenging even for fine-tuned LMs, which means predicting exactly how many authors have contributed to a text and exactly who wrote each span of text are extremely hard in this multi-LLM scenario, irrespective of how few or how many LLMs are collaborating.
6 Conclusion
We present CollabStory, the first exclusively LLM-LLM or machine-machine collaborative story dataset, and demonstrate the tasks it enables. We find that LLMs are able to collectively generate creative stories at par with human-written stories via sequential prompting. Using this dataset, we demonstrate which multi-LLM authorship tasks are most challenging. Recent developments have significantly advanced LLM-assisted writing, sparking widespread discussions about the nature of authorship. Beyond using LLMs for paraphrasing, editing, and enhancing text, there exists an extreme scenario where text is generated entirely by multiple LLMs. Our work addresses this extreme case, raising several nuanced authorship concerns: Who should be considered the true creative source in such a situation? Should all LLMs involved be credited? Or should the human developers designing the prompts be acknowledged as the primary authors? Moreover, should the LLM that contributed the most—whether in terms of word count, narrative depth, or plot twists—be granted greater ownership? We will soon need “Catch As Catch Can” methods to not only find all points where authorship changes within an article but also simultaneously attribute each independent segment to the specific LLM author. As more and more LLMs are becoming easier to access, malicious actors could combine texts from different LLMs to evade automated and in-built misinformation flaggers, or students might circumvent credibility checks by having different LLMs write different sections of an academic article. We further discuss the real-world implications of such tasks in the Appendix (Section B). Thus, CollabStory has been developed as a resource with long-form stories written by multiple LLMs to support the development and expansion of tasks and methods that can help address incoming challenges brought by LLM-LLM interactions.
Limitations
Ethics Statement
Appendix A Appendix
A.1 LLM Prompting
A.2 Dataset Cleaning
Feature Author K=1 K=2 K=3 K=4 K=5
Llama 0.60 ± 0.05 0.61 ± 0.05 0.59 ± 0.05 0.58 ± 0.05 0.58 ± 0.06
mistral 0.64 ± 0.04 0.62 ± 0.04 0.61 ± 0.05 0.61 ± 0.04 0.61 ± 0.05
Olmo 0.62 ± 0.05 0.57 ± 0.06 0.58 ± 0.06 0.58 ± 0.06 0.57 ± 0.07
Orca 0.62 ± 0.05 0.62 ± 0.04 0.61 ± 0.04 0.61 ± 0.05 0.60 ± 0.05
Llama 83.11 ± 8.35 82.61 ± 8.50 82.75 ± 9.00 80.88 ± 10.10 80.13 ± 9.19
Mistral 81.04 ± 8.58 83.99 ± 8.40 81.59 ± 10.07 82.02 ± 9.09 79.91 ± 9.44
Olmo 80.78 ± 9.01 83.31 ± 9.87 81.41 ± 9.81 80.55 ± 10.73 80.45 ± 10.20
Orca 83.08 ± 8.54 82.51 ± 8.45 80.86 ± 9.65 79.95 ± 9.97 79.51 ± 9.80
Coherence Gemma 0.49 ± 0.07 0.47 ± 0.08 0.47 ± 0.08 0.47 ± 0.08 0.48 ± 0.08
Llama 0.44 ± 0.08 0.47 ± 0.08 0.47 ± 0.08 0.47 ± 0.08 0.49 ± 0.08
mistral 0.46 ± 0.07 0.44 ± 0.07 0.45 ± 0.07 0.45 ± 0.07 0.46 ± 0.07
Olmo 0.47 ± 0.08 0.43 ± 0.07 0.45 ± 0.08 0.45 ± 0.08 0.45 ± 0.08
Orca 0.44 ± 0.07 0.45 ± 0.08 0.45 ± 0.08 0.46 ± 0.08 0.47 ± 0.07
@@ A.2 Dataset Cleaning
Appendix B Authorship Analysis: Real world implications
This work have profound implications for various stakeholders in the burgeoning socio-technical system of generative AI. Our research introduces authorship-related tasks using CollabStory, which can help address these concerns by accurately discerning the usage of multiple LLMs in texts (summarized in Table 6. Our extension of PAN-inspired authorship tasks is closely linked to real-world implications, as follows: